### v2

# DISL: Fueling Research with A Large Dataset of Solidity Smart Contracts

Abstract
The DISL dataset features a collection of 514,506514506514,506514 , 506 unique Solidity files that have been deployed to Ethereum mainnet. It caters to the need for a large and diverse dataset of real-world smart contracts. DISL serves as a resource for developing machine learning systems and for benchmarking software engineering tools designed for smart contracts. By aggregating every verified smart contract from Etherscan up to January 15, 2024, DISL surpasses existing datasets in size and recency.
1 Introduction
Smart contracts are self-executing programs running on top of a decentralized ledger [23]. Their use-cases range from financial applications, governance, gaming to art. Solidity stands as one of the most used programming languages for smart contract development.
Security and safety of smart contracts is crucial because their deployment on an immutable ledger makes vulnerabilities and bugs irrevocable, risking significant financial losses. Thus, it is essential to conduct thorough testing, analysis, and repair of buggy smart contract code. Tools and techniques employed for this purpose use both static to dynamic analysis [16, 19]. The performance of these tools however is overshadowed by the number of recent successful attacks performed on real projects deployed on chain [11]. Recent studies indicate that the reliance on synthetic, small, and non-diverse smart contract datasets fail to represent the complexity of real-world projects. Those small datasets hinder the development and assessment of reliable tools, a core reason for their ineffectiveness [17, 21, 16].
A dataset of real-world smart contracts will aid researchers and developers in different tasks, including the production of smart contract development tools. Moreover, as AI-assisted tools gain popularity in smart contract analysis, repair, and synthesis tasks, the demand for large datasets for sake of machine learning tasks on contracts grows increasingly critical [14, 24, 13].
Yet, we notice that the current datasets of Solidity smart contracts are limited: they are either outdated (old Solidity versions) and do not reflect real-world deployed contracts (See section 5 for a discussion).
To address this problem, we collected DISL, a large dataset of real-world smart contracts. This dataset is useful: 1) For assessing smart contract analysis tools 2) For training machine learning and LLM-based tools for smart contracts.
A verified smart contract on Etherscan is a contract whose source code has been uploaded to Etherscan and successfully matched with its compiled code on blockchain to provide transparency and public accessibility to the source.
DISL only includes the verified source code of smart contracts to ensure it comprises solely real contracts in use. DISL contains the source code for all smart contracts on Etherscan from the genesis block to January 15, 2024. DISL is processed with a deduplication phase in order to remove the Solidity smart contracts that appear thousands of times on Ethereum.
To sum up our contributions are:
  • The largest dataset of Ethereum smart contracts, containing full source code and metadata satisfying the requirements of a suitable dataset for research and tool development (see 2.2). Additionally, as the focus of related datasets (Section 5) is Solidity contracts, DISL (raw collection) with 7,18871887,1887 , 188 smart contracts in the Vyper programming language, is the largest dataset of Vyper contracts currently available.
  • A publicly-available version published on Huggingface accessible for researchers and practitioners: https://huggingface.co/datasets/ASSERT-KTH/DISL.
2 Dataset Collection Process
2.1 Definitions
Here, we clarify the terminology.
  • Deployed contract: A smart contract address on Ethereum blockchain main network that is associated with a binary code stored in an Ethereum block.
  • Raw contract: A concatenated version of all source codes (Solidity files) retrieved from Etherscan for a deployed contract. It may contain multiple libraries used within the given contract address.
  • Solidity file: The source code files that is used to write Solidity smart contract code. A raw contract may contain several Solidity files, which themselves may contain more than one “contract” keyword.
2.2 Requirements
We want to collect a dataset of smart contracts that meets the following requirements.
  • The dataset shall contain smart contracts covering a range of different applications (DeFi, art, etc.)
  • The dataset shall contain Solidity smart contracts written in recent versions of Solidity (including 2022 onwards).
  • The dataset shall be suitable for AI tasks, hence should contain limited duplication.
  • The dataset shall be available in a widely-usable format, supported by mainstream dataset platforms.
2.3 Initial dataset.
We start our work with the Andstor dataset, which includes every deployed Ethereum smart contract up to April 1, 2022 [10]. This dataset contained a total of 2,217,69222176922,217,6922 , 217 , 692 entries of raw contracts and 186,397186397186,397186 , 397 lines of deduplicated Solidity files. DISL extends Andstor.
2.4 Collection of most recent contracts.
Next, we use the Google BigQuery database for Ethereum that provides access to blockchain data for analysis [1]. This daily updated database allows exploration of smart contract transactions. We collected smart contracts data starting from the last day of the previously mentioned dataset, April 1, 2022, for all deployed contracts that at least have one transaction.
1 SELECT contracts.address, COUNT(1) AS tx_count
2 FROM bigquery-public-data.crypto_ethereum.contracts AS contracts
3 JOIN bigquery-public-data.crypto_ethereum.transactions AS transactions
4 ON (transactions.to_address = contracts.address)
5 WHERE transactions.block_timestamp >= TIMESTAMP("2022-04-01")
6 GROUP BY contracts.address
7 ORDER BY tx_count DESC;
Listing 1: Query used for retrieving the dataset
By using Google BigQuery we collected a CSV file with 2,709,03027090302,709,0302 , 709 , 030 entries, where every line has the address of the deployed contract and the number of transactions.
Using Etherscan’s public APIs, we retrieved the source code for all listed addresses from the data collected on Google BigQuery, successfully acquiring data for 2,660,65826606582,660,6582 , 660 , 658 contracts stored in JSON format . Next, we remove the rows composed of empty JSON files, which indicates when a contract does not have a verified address on Etherscan. This results in what we call a raw dataset, consisting of 1,080,57910805791,080,5791 , 080 , 579 rows.
We merged our parquet files with Andstor dataset to obtain 3,298,27132982713,298,2713 , 298 , 271 rows, each one corresponding to a deployed contract.
2.5 Deduplication
The structure of deployed contracts is usually a mix of dependencies and project-specific contract definitions that inherit or use the dependencies. The dependencies employed in deployed contracts originate from popular smart contract libraries, including OpenZeppelin, Safe, and Provable [3, 4, 7]. This means source codes in DISL raw dataset may contain many times the same Solidity library code. Therefore, we perform deduplication (only for decomposed collection) to ensure that all dataset entries have unique value.
[…]
Table 1: Details of smart contract metadata in each of the two raw and decomposed collections of the dataset (each row is a columns in the dataset table)
Column Description Raw Decomposed
Contract name Contract name
Contract address Address of the contract on the Ethereum blockchain
Language Language of the contract (Solidity, Vyper)
Source code In raw, it contains all the code provided by the Etherscan API appended together, in decomposed the source code of a single Solidity file
Compiler version Version of the compiler used to compile the smart contract
License type Name of the license of the smart contract
ABI Application Binary Interface of the contract $\times$×
Optimization used Boolean for running optimization $\times$×
Runs Number of runs in optimization $\times$×
Constructor arguments Arguments for the constructor of the smart contract $\times$×
EVM version Version of the Ethereum Virtual Machine used to compile the contract $\times$×
Library Name and address of the library used to compile the contract $\times$×
Proxy Boolean true if the contract is a proxy $\times$×
Implementation Address of the implementation if the contract is a proxy $\times$×
Swarm source Address of the source code in Swarm $\times$×
File path In the decomposed dataset, contains the path to the file in the contract structure $\times$×
3 Dataset Content
As mentioned in Section 2, DISL consists of two collections raw and decomposed. Table 2 provides the overview of DISL. It contains more than three millions of deployed smart contracts. The major outcome of this work is a dataset of 514,506514506514,506514 , 506 real-worl smart contracts (decomposed collection), with no duplication (last row of Table 2). We publish both of these collections as tabular data in Huggingface (https://huggingface.co/datasets/ASSERT-KTH/DISL).
Table 2: Overview of DISL dataset collections
Raw
# Deployed smart contracts 3,298,271
Solidity file count (decomposed) 12,931,943
Solidity file count (decomposed, deduplicated) 514,506
File Content: Each collection is divided into several tabular files to improve manageability. Every file is stored in Parquet format and contains up to 30,0003000030,00030 , 000 rows.
Metadata: We add metadata provided by Etherscan as columns in the tabular format of the dataset. Table 1 shows the available metadata.
Binary: The dataset contains all the metadata to actually reconstruct the bytecode representation via compilation. The deployed bytecode of the contract on blockchain is also available with the collected contract addresses.
4 Applications
Thanks to its size and nature (only verified smart contracts), the DISL dataset offers significant advantages in two primary areas: AI-based tool development and benchmarking of smart contract software engineering tools.
Task-specific training of large language models (LLM) is the process of adapting LLM parameters to improve performance on specific tasks (e.g. code synthesis). Fine-tuning is a proven technique for enhancing the performance of LLMs on code-related tasks [12, 18]. DISL is a valuable candidate for such fine-tuning operations as it contains a large smart contract corpora with deduplicated files [22]
Benchmarking involves evaluating the performance of software engineering tools (traditional and AI-based) using a standard dataset to measure their performance. As DISL contains unique, real-world contracts, it is a new valuable benchmarking suite compared to the existing ones.
Empirical studies require real scenarios. However, academic smart contract analysis tools lack real-life examples [24, 17]. The DISL dataset serves as a valuable resource for studying contracts.
5 Related Work
Table 3: Related smart contract source datasets and their size (number of source files)
Dataset Year Size
Andstor (deduplicated[10] 2022 186,397186397186,397186 , 397
Fiesta (deduplicated) [^1] It uses hashes for deduplication. [6] 2023 149,386149386149,386149 , 386
Sanctuary [20] 2022 144,857144857144,857144 , 857
SmartBugs-Wild [8, 15] 2020 47,5874758747,58747 , 587
Ren et al.  [21, 5] 2021 46,1864618646,18646 , 186
DAppSCAN [2] 2023 39,9043990439,90439 , 904
DISL (raw) 2024 3,298,27132982713,298,2713 , 298 , 271
DISL (deduplicated) 2024 514,506514506514,506514 , 506
1. It uses hashes for deduplication.
Table 3 summarizes the related work on datasets of smart contracts.
Andstor contains 186,397186397186,397186 , 397 deduplicated Solidity files [10]. This dataset served as the initial seed for our work. This dataset uses the same deduplication method as DISL.
The Fiesta dataset contains 149,386149386149,386149 , 386 contracts [6]. It uses hashing to deduplicate contracts.
Sanctuary has 149,386149386149,386149 , 386 contracts from Ethereum mainnet, among them 71,4947149471,49471 , 494 are unique. According to their website, it has been updated for the last time on the 24th of January 2022. It also contains contracts from testnets of Ethereum and is part of a project that includes other networks. [20]
SmartBugs-Wild and Meng Ren’s dataset, containing 47,5874758747,58747 , 587 and 46,1864618646,18646 , 186 unique smart contracts respectively, have been used in vulnerability analysis and tool evaluation [8, 5].
DAppSCAN contains 39,904 audited smart contracts. DAppSCAN’s relatively small size and focus on audited projects limits its utility for machine learning tasks [2].
DISL is the largest dataset of smart contracts to date, with over $3.73.7$ million raw smart contract records and a unique subset of 514,506514506514,506514 , 506 deduplicated Solidity files.
6 Conclusion
DISL is the largest dataset of smart contract source files at the time of writing. It includes 3,298,271 raw and 514,506 deduplicated Solidity source files, all taken from real-world, deployed, verified contracts deployed on Ethereum. We envision that DISL will fuel future research on smart contracts.